5. Supply costs at Lattea Corporation's chain of gyms are listed below:

Client-Visits Supply Cost
  March 11,670      $28,584     
  April 11,466      $28,418     
  May 11,998      $28,842     
  June 14,300      $28,938     
  July 11,730      $28,645     
  August 11,216      $28,244     
  September 12,010      $28,843     
  October 11,701      $28,601     
  November 11,849      $28,726     

Management believes that supply cost is a mixed cost that depends on client-visits. Use the high-low method to estimate the variable and fixed components of this cost, Compute the variable component first, rounding off to the nearest whole cent. Then compute the fixed component, rounding off to the nearest whole dollar. Those estimates would be closest to: (Round your Variable cost per unit to 2 decimal places.)
Answer #1

Variable cost per visit = Change in Cost/Change in visit = (28938-28244)/(14300-11216) = 0.23

Fixed cost = Total Cost-Variable cost = 28938-(14300*.23) = 25649